OpenStudy (anonymous):

Help! The annual salaries of all employees at a financial company are normally distributed with a mean= $34,000 and a standard deviation= $4,000. What is the z-score of a company employee who makes an annual salary of $28,000?

OpenStudy (anonymous):

Wouldn't it be –1.5?

OpenStudy (anonymous):

-1.5

OpenStudy (anonymous):

how do you get -1.5 ?

OpenStudy (anonymous):

$28,000 is $6,000 off from $34,000. $4,000(1.5) = $6,000 and since $28,000 is less than $34,000, the 1.5 is negative =-1.5

OpenStudy (anonymous):

uhm.. what do you understand by z-score

OpenStudy (anonymous):

I understand that I learned about it MONTHS ago, and apparently don't remember anything about it. XD

OpenStudy (rational):

-1.5 is right

OpenStudy (rational):

zscore = (observation - mean)/stddev = (28k - 34k)/4k = -6/4 = -1.5

OpenStudy (anonymous):

Yay! @rational You wouldn't happen to willing to help me on another problem like this would you?

OpenStudy (rational):

wil try, ask...

OpenStudy (anonymous):

A simple random sample of 50 adults was surveyed, and it was found that the mean amount of time that they spend surfing the Internet per day is 54.2 minutes, with a standard deviation of 14.0 minutes. What is the 99% confidence interval for the number of minutes that an adult spends surfing the Internet per day? 49.1 minutes to 59.3 minutes 50.3 minutes to 58.1 minutes 54.2 minutes to 58.1 minutes 54.2 minutes to 59.3 minutes

OpenStudy (anonymous):

please?

OpenStudy (anonymous):

@rational ^^^^